OpenStudy (anonymous):

give that a=1/4f , using the quadratic equation y=2x^2, determine the value of f. Can someone help?

OpenStudy (triciaal):

in the quadratic equation y = 2 x^2 a = 2 given that a = 1/4 f then 1/4 f = 2 multiply each side of the equation by 4 gives f = 8